# Characterizing graph-nonedge pairs with single interval Cayley configuration spaces in 3-dimensions

### v1

@@ C.3.4. Proof of Proposition 8 [Minimal pair inductive structure]
Lemma 46 (Additional properties of graphs with $H^{-1}$ in Figures 15(a) - 15(m)).
Let $G$ be a graph with subgraphs $G_{1}$, $G_{2}$, and $S$ such that $G=G_{1}\cup G_{2}$, $G_{1}$ is one of the graphs in Figures 15(a) - 15(m), $S=G_{1}\cap G_{2}$ is the subgraph of $G_{1}$ induced by some subset of $\{w_{i}\}$ that contains some nonedge $e$, and $G\setminus S$ is disconnected. Also, let $e=w_{1}w_{3}$ if $G_{1}$ is the graph in Figure 15(c). If $G_{2}\cup e$ has an $e$-preserving forbidden minor, then the red edge of $G_{1}$ is preserved in some forbidden minor of $G$.
@@
Assume that $G_{2}\cup e$ has an $e$-preserving forbidden minor. We will show that $G$ has an $f$-preserving forbidden minor, where $f$ is the red edge of $G_{1}$. If there exists a minor $[G]$ in which $G_{2}$ is preserved and $[f]=[e]$, then this follows from the fact that $G_{2}\cup e$ has an $e$-preserving forbidden minor. Using this fact, it is easy to see that the lemma is true if either (i) $G_{1}$ is one of the graphs in Figures 15(e), 15(f), 15(k), 15(l), or 15(m), (ii) $e=w_{1}w_{3}$ and $G_{1}$ is one of the graphs in Figures 15(d) or 15(i), or (iii) $e=w_{1}w_{4}$ and $G_{1}$ is one of the graphs in Figures 15(g) or 15(h).
Next, consider the case where $G_{1}$ is the graph in Figure 15(a). Note that either $w_{1}w_{2}$ or $w_{2}w_{3}$ is preserved in any $e$-preserving minor of $G_{2}$. Also, observe that there exists a minor $[G]$ in which $G_{2}$ is preserved, $[e]$ is an edge, and $[f]$ is either $[w_{1}w_{2}]$ or $[w_{2}w_{3}]$. Combining these facts shows that $G$ has an $f$-preserving forbidden minor. A similar argument applies if either (i) $e=w_{1}w_{3}$ and $G_{1}$ is one of the graphs in Figures 15(b), 15(c), 15(g), 15(h), or 15(j) or (ii) $e=w_{2}w_{4}$ and $G_{1}$ is one of the graphs in Figures 15(d), 15(i), or 15(j).
@@
Lemma 47 (Additional structure necessary for induction).
Let $(G,f)$ be a minimal graph-nonedge pair with a top-level expanded $f$-separating cms $C^{-1}$ whose expanded $f$-component is $H^{-1}$, and let $G^{\prime}=G\setminus(H^{-1}\setminus C^{-1})$. There exists a nonedge $f^{\prime}$ of $C^{-1}$ such that some atom of $G^{\prime}\cup f^{\prime}$ contains $f^{\prime}$ and has an $f^{\prime}$-preserving forbidden minor and the pair $(G^{\prime},f^{\prime})$ has no reducing edge. Furthermore, these statements are true when (i) $f^{\prime}=w_{1}w_{3}$ and $H^{-1}$ is the graph in Figure 15(c), $f$ is the red edge, and $C^{-1}$ is the subgraph induced by $\{w_{i}\}$, or (ii) $f^{\prime}$ is any nonedge of $C^{-1}$ and $H^{-1}$ is one of the graphs in Figures 15(d), 15(i), or 15(j), $f$ is the red edge, and $C^{-1}$ is the subgraph induced by $\{w_{i}\}$.
Proof.
@@
Since any choice for $f^{\prime}$ that satisfies (a) does not satisfy (b), the pair $(G^{\prime},f^{\prime})$ has some reducing edge $h$. Consider the minor $[G\cup f]$ obtained by contracting $h$. By definition, $f^{\prime}$ is retained in $[G^{\prime}]$ and some induced subgraph $K$ of $[G^{\prime}]$ is such that $K\cup[f^{\prime}]$ is an atom that contains $[f^{\prime}]$ and has an $[f^{\prime}]$-preserving forbidden minor. Let $S=[C^{-1}]$, $e=[f^{\prime}]$, $G_{1}=[H^{-1}]$, and $G_{2}=K$. By Lemma 44, $H^{-1}$ is one of the graphs in Figures 15(a) - 15(m) and $C^{-1}$ is the subgraph induced by the vertex set $\{w_{i}\}$. There are two cases.
@@
Since $h$ is contained in $G^{\prime}$, it must be contained in $S$. Combining this with the fact that $e$ is retained in $[G\cup f]$ shows that $H^{-1}$ is one of the graphs in Figures 15(b), 15(f), 15(g), 15(h), 15(k), 15(l), or 15(m). There are $9$ subcases.
@@
Subcase 2: $H^{-1}$ is one of the graphs in Figures 15(f), 15(k), 15(l), or 15(m).
@@
Subcase 3: $H^{-1}$ is one of the graphs in Figures 15(g) or 15(h), $uw_{1}$ is an edge, and $h$ is $w_{2}w_{3}$ or $w_{2}w_{4}$.
@@
Subcase 4: $H^{-1}$ is one of the graphs in Figures 15(g) or 15(h), $uw_{1}$ is an edge, $h$ is $w_{3}w_{4}$, and $f^{\prime}=w_{1}w_{3}$.
Let $u$ be the endpoint of $f$ not contained in $C^{-1}$. We will show that $uw_{1}$ is reducing with respect to $(G,f)$. Let $[G\cup f]_{A}$ be the minor obtained by contracting $uw_{1}$. Observe that $f$ is retained in $[G\cup f]_{A}$, and so $uw_{1}$ is not of Type (3). We show that $uw_{1}$ is not of Type (1) or (2) by identifying some atom of $[G\cup f]_{A}$ that contains $[f]_{A}$ and has an $[f]_{A}$-preserving forbidden minor. It suffices to show that some atom of $G^{\prime}\cup\{w_{1}w_{3},w_{1}w_{4}\}$ contains $w_{1}w_{4}$ and has a $w_{1}w_{4}$-preserving forbidden minor.
@@
Subcase 5: $H^{-1}$ is one of the graphs in Figures 15(g) or 15(h), $uw_{1}$ is an edge, $h$ is $w_{3}w_{4}$, and $f^{\prime}=w_{1}w_{4}$.
This case is similar to Subcase 4. We will show that $uw_{1}$ is reducing with respect to $(G,f)$. Let $[G\cup f]_{A}$ be the minor obtained by contracting $uw_{1}$. Observe that $f$ is retained in $[G\cup f]_{A}$, and so $uw_{1}$ is not of Type (3). We show that $uw_{1}$ is not of Type (1) or (2) by identifying some atom of $[G\cup f]_{A}$ that contains $[f]_{A}$ and has an $[f]_{A}$-preserving forbidden minor. It suffices to show that some atom of $G^{\prime}\cup\{w_{1}w_{3},w_{1}w_{4}\}$ contains $w_{1}w_{4}$ and has a $w_{1}w_{4}$-preserving forbidden minor.
@@
Let $x$ be any vertex of $H^{-1}\setminus C^{-1}$ other that an endpoint of $f$. We will show that $xw_{1}$ is reducing with respect to $(G,f)$. Observe that the minor $[G\cup f]_{A}$ obtained by contracting $xw_{1}$ is $f$-retaining, and so $xw_{1}$ is not of Type (3). We show that $xw_{1}$ is not of Type (1) or (2) by identifying some atom of $[G\cup f]_{A}$ that contains $[f]_{A}$ and has an $[f]_{A}$-preserving forbidden minor.
Since $f^{\prime}$ satisfies (a), some atom $J$ of $G^{\prime}\cup f^{\prime}$ contains $f^{\prime}$ and has an $f^{\prime}$-preserving forbidden minor. We will show that either $J$ contains $w_{3}$ or $w_{2}w_{3}$ is reducing with respect to $(G^{\prime},f^{\prime})$. In the latter case, Subcase 6 applies. If $J$ does not contain $w_{3}$, then note that the minor $[G\cup f^{\prime}]_{B}$ obtained by contracting $w_{2}w_{3}$ is $f^{\prime}$-retaining, so it $w_{2}w_{3}$ is not of Type (3). Also, some cms $E$ contained in $J$ is such that all paths in $G^{\prime}\cup f^{\prime}$ between $w_{3}$ and any vertex of $J\setminus E$ contain some vertex of $E$. Hence, $J$ is weakly retained in $[G^{\prime}\cup f^{\prime}]_{B}$. Therefore, some atom of $[G^{\prime}\cup f^{\prime}]_{B}$ contains $[J]_{B}$, and consequently contains $[f^{\prime}]_{B}$ and has an $[f^{\prime}]_{B}$-preserving forbidden minor. This shows that $w_{2}w_{3}$ is not of Type (1) or (2), and so it is reducing with respect to $(G^{\prime},f^{\prime})$.
Next, since $J$ contains $w_{3}$, $J$ is some induced subgraph $G_{2}$ of $[G^{\prime}]_{A}$ such that $f^{\prime}$ is $[f^{\prime}]_{A}$. Let $G_{1}=[C^{-1}]_{A}\cup[u]_{A}$, $S=G_{1}\cap G_{2}$, and $e=[w_{1}w_{3}]_{A}$. Then, $G_{1}\cup G_{2}$ contains $[f]_{A}$. Also, since $G_{2}$ has an $[f^{\prime}]_{A}$-preserving forbidden minor, it is easy to see that $G_{1}\cup G_{2}$ has an $[f]_{A}$-preserving forbidden minor. Furthermore, since $G_{1}$ and $G_{2}$ are atoms and $e$ is a nonedge, Lemma 64 shows that $G_{1}\cup G_{2}$ is an atom. Furthermore, $G_{1}\cup G_{2}$ is contained in some atom of $[G\cup f]_{A}$ since it is an induced subgraph of $[G\cup f]_{A}$, and this atom has the desired properties.
@@
Finally, by the above arguments, some induced subgraph $I$ of $[G^{\prime}]_{A}$ is such that $I\cup[w_{1}w_{3}]_{A}$ is an atom that contains $[w_{1}]_{A}$, $[w_{3}]_{A}$, and $[w_{4}]_{A}$ and has a $[w_{1}w_{4}]_{A}$-preserving forbidden minor. Let $G_{1}=[C^{-1}]_{A}\cup[u]_{A}$, $G_{2}=I$, $S=G_{1}\cap G_{2}$, and $e=[w_{1}w_{3}]_{A}$. Similar to the argument in Subcase 8, we can use Lemma 64 to show that some atom of $[G\cup f]_{A}$ that contains $[f]_{A}$ and has an $[f]_{A}$-preserving forbidden minor.
@@
Finally, we prove Lemma 48, below, and then prove Proposition 8.
Lemma 48 (Structure of $G^{\prime}\cup f^{\prime}$).
@@
Assume to the contrary that $G^{\prime}\cup f^{\prime}$ cannot be obtained from $J$ by adding vertices whose neighborhood are $\{u,v\}$. Then, there exists a cms $E$ of $G^{\prime}\cup f^{\prime}$ contained in $J$ such that $E$ contains some vertex $x$ that is either the only vertex in $E$ or some vertex other than $u$ or $v$. Also, there exists a vertex $y$ not contained in $J$ and such that $xy$ is an edge of $G^{\prime}\cup f^{\prime}$ and all paths in $G^{\prime}\cup f^{\prime}$ between $y$ and any vertex in $J\setminus E$ contains some vertex in $E$. Hence, $J$ is weakly retained in the minor $[G^{\prime}\cup f^{\prime}]$ obtained by contracting $xy$. Furthermore, $[G^{\prime}\cup f^{\prime}]$ is $f^{\prime}$-retaining and contains some atom that contains $[f^{\prime}]$ and has an $[f^{\prime}]$-preserving forbidden minor. The above facts show that $xy$ is reducing with respect to $(G^{\prime},f^{\prime})$, which is a contradiction. This completes the proof. ∎
Proof of Proposition 8.
The proposition follows easily from Lemmas 47 and 48. ∎
C.4. Proof of Proposition 9 [Pruned possible graphs for top-level $H^{-1}$]
Lemma 49 (Vertices in atom containing $f^{\prime}$ 1).
Consider a minimal graph-nonedge pair $(G,f)$ with a top-level expanded $f$-separating cms $C^{-1}$ whose expanded $f$-component is $H^{-1}$, and let $G^{\prime}=G\setminus(H^{-1}\setminus C^{-1})$. Also, let $C^{-1}$ be the graph in Figure 13(b). For any nonedge $f^{\prime}$ of $C^{-1}$, if some atom $J$ of $G^{\prime}\cup f^{\prime}$ contains $f^{\prime}$ and has an $f^{\prime}$-preserving forbidden minor and the pair $(G^{\prime},f^{\prime})$ has no reducing edge, then $J$ contains $w_{1}$, $w_{3}$, and $w_{4}$.
Proof.
@@ C.4. Proof of Proposition 9 [Pruned possible graphs for top-level $H^{-1}$]
Wlog, assume $f^{\prime}=w_{1}w_{3}$. Then, $J$ contains $w_{1}$ and $w_{3}$. If $J$ does not contain $w_{4}$, then Lemma 48 states that the neighborhood of $w_{4}$ in $G^{\prime}$ is $\{w_{1},w_{3}\}$. However, Lemma 4 Statement (2) can be used to show that $w_{4}$ neighbors some vertex of $G^{\prime}$ not in $\{w_{1},w_{3}\}$. This is a contradiction, and thus $J$ contains $w_{4}$. ∎
Lemma 50 (Vertices in atom containing $f^{\prime}$ 2).
Consider a minimal graph-nonedge pair $(G,f)$ with a top-level expanded $f$-separating cms $C^{-1}$ whose expanded $f$-component is $H^{-1}$, and let $G^{\prime}=G\setminus(H^{-1}\setminus C^{-1})$. Also, let $C^{-1}$ be the graph in Figure 13(d). For any nonedge $f^{\prime}$ of $C^{-1}$, if some atom $J$ of $G^{\prime}\cup f^{\prime}$ contains $f^{\prime}$ and has an $f^{\prime}$-preserving forbidden minor and the pair $(G^{\prime},f^{\prime})$ has no reducing edge, then $J$ contains $C^{-1}$.
Proof.
Wlog, assume that $f^{\prime}=w_{1}w_{3}$. By assumption, $J$ contains $w_{1}$ and $w_{3}$. If $J$ does not contain some vertex in $\{w_{2},w_{4}\}$, then Lemma 48 shows that the neighborhood of this vertex in $G^{\prime}$ is $\{w_{1},w_{3}\}$. However, observe that $w_{2}w_{4}$ is an edge of $G^{\prime}$. Thus, $J$ contains $C^{-1}$. ∎
Lemma 51.
Additional pruning of cases for $H^{-1}$ For any minimal graph-nonedge pair $(G,f)$ with a top-level expanded $f$-separating cms $C^{-1}$, the expanded $f$-component of $C^{-1}$ is one of the graphs in Figures 15(a) - 15(h), where $C^{-1}$ is the subgraph induced by the vertex set $\{w_{i}\}$. Furthermore, if $H^{-1}$ is one of the graphs in Figures 15(c), 15(d), or 15(e), then the dashed line-segment is a nonedge.
Proof.
@@
By Lemma 44, the expanded $f$-component $H^{-1}$ of $C^{-1}$ is one of the graphs in Figures 15(a) - 15(m), where $C^{-1}$ is the subgraph induced by $\{w_{i}\}$. Furthermore, if $H^{-1}$ is one of the graphs in Figures 15(c), 15(d), 15(e), 15(i), 15(k), or 15(l), then at least one dashed line-segment is a nonedge. For each case of $H^{-1}$ in Figures 15(i) - 15(m), we will show that $(G,f)$ has a reducing edge, which contradicts the minimality of this pair.
Case 1: $H^{-1}$ is one of the graphs Figures 15(i) or 15(j).
@@
We prove the lemma in the case where $H^{-1}$ is the graph Figure 15(i). The proof in the case where $H^{-1}$ is the graph Figure 15(j) is almost identical, so we omit it. Let $G^{\prime}=G\setminus(H^{-1}\setminus C^{-1})$. By Lemma 47, some atom $J$ of $G^{\prime}\cup w_{1}w_{3}$ contains $w_{1}w_{3}$ and has a $w_{1}w_{3}$-preserving forbidden minor and the pair $(G^{\prime},w_{1}w_{3})$ has no reducing edge. Also, Lemma 49 shows that $J$ contains $w_{4}$. If $J$ contains $w_{2}$, then let $x$ be any vertex of $H^{-1}\setminus C^{-1}$ that neighbors $w_{3}$. We will show that $xw_{3}$ is reducing with respect to $(G,f)$. Note that $f$ is retained in the minor $[G\cup f]$ obtained by contracting $xw_{3}$, and so this edge is not of Type (3). To see that $xw_{3}$ is not of Type (1) or (2), observe that $J$ is some induced subgraph $G_{2}$ of $[G^{\prime}]$ such that $w_{1}w_{3}$ is $[w_{1}x]$. Hence, $G_{2}$ is an atom that contains $S=[C^{-1}]$ and has a $[w_{1}x]$-preserving forbidden minor. Let $G_{1}$ be the $\{[u],[w_{1}],[w_{3}]\}$-component of $[H^{-1}]$ that contains $S$. Then, $G_{1}$ is an atom and $S=G_{1}\cap G_{2}$. Since $S$ contains the nonedge $[w_{2}w_{4}]$, Lemma 64 shows that $G_{1}\cup G_{2}$ is an atom. Lastly, $G_{1}\cup G_{2}$ clearly has an $[f]$-preserving forbidden minor. Therefore, the atom of $[G\cup f]$ that contains $G_{1}\cup G_{2}$ contains $[f]$ and has an $[f]$-preserving forbidden minor.
@@
Finally, observe that $K$ is some induced subgraph $G_{1}$ of $[G^{\prime}]$ such that $w_{2}w_{4}$ is $[xw_{4}]$. Let $G_{2}=[H^{-1}]$ and observe that $S=G_{1}\cap G_{2}=[C^{-1}]$ contains the nonedge $[w_{1}w_{3}]$. Since $G_{1}$ and $G_{2}$ are atoms, Lemma 64 shows that $G_{1}\cup G_{2}$ is an atom. Observe that $G_{1}\cup G_{2}$ has an $[f]$-preserving forbidden minor. These facts show that the atom of $[G\cup f]$ that contains $G_{1}\cup G_{2}$ contains $[f]$ and has an $[f]$-preserving forbidden minor. Therefore, $xw_{2}$ is not of Type (1) or (2), as desired.
Case 2: $H^{-1}$ is one of the graphs in Figures 15(k), 15(l), or 15(m).

### v2

@@ B.2.2   Proof of Lemma 33
Lemma 55 (Additional properties of graphs with $H^{-1}$ in Figures 20(a) - 20(m)).
Let $G$ be a graph with subgraphs $G_{1}$, $G_{2}$, and $S$ such that $G=G_{1}\cup G_{2}$, $G_{1}$ is one of the graphs in Figures 20(a) - 20(m), $S=G_{1}\cap G_{2}$ is the subgraph of $G_{1}$ induced by some subset of $\{w_{i}\}$ that contains some nonedge $e$, and $G\setminus S$ is disconnected. Also, let $e=w_{1}w_{3}$ if $G_{1}$ is the graph in Figure 20(c). If $G_{2}\cup e$ has an $e$-preserving forbidden minor, then the red edge of $G_{1}$ is preserved in some forbidden minor of $G$.
@@
Assume that $G_{2}\cup e$ has an $e$-preserving forbidden minor. We will show that $G$ has an $f$-preserving forbidden minor, where $f$ is the red edge of $G_{1}$. If there exists a minor $[G]$ in which $G_{2}$ is preserved and $[f]=[e]$, then this follows from the fact that $G_{2}\cup e$ has an $e$-preserving forbidden minor. Using this fact, it is easy to see that the lemma is true if either (i) $G_{1}$ is one of the graphs in Figures 20(e), 20(f), 20(k), 20(l), or 20(m), (ii) $e=w_{1}w_{3}$ and $G_{1}$ is one of the graphs in Figures 20(d) or 20(i), or (iii) $e=w_{1}w_{4}$ and $G_{1}$ is one of the graphs in Figures 20(g) or 20(h).
Next, consider the case where $G_{1}$ is the graph in Figure 20(a). Note that either $w_{1}w_{2}$ or $w_{2}w_{3}$ is preserved in any $e$-preserving minor of $G_{2}$. Also, observe that there exists a minor $[G]$ in which $G_{2}$ is preserved, $[e]$ is an edge, and $[f]$ is either $[w_{1}w_{2}]$ or $[w_{2}w_{3}]$. Combining these facts shows that $G$ has an $f$-preserving forbidden minor. A similar argument applies if either (i) $e=w_{1}w_{3}$ and $G_{1}$ is one of the graphs in Figures 20(b), 20(c), 20(g), 20(h), or 20(j) or (ii) $e=w_{2}w_{4}$ and $G_{1}$ is one of the graphs in Figures 20(d), 20(i), or 20(j).
@@
Proof of Lemma 33.
@@
Since any choice for $f^{\prime}$ that satisfies (a) does not satisfy (b), the pair $(G^{\prime},f^{\prime})$ has some reducing edge $h$. Consider the minor $[G\cup f]$ obtained by contracting $h$. By definition, $f^{\prime}$ is retained in $[G^{\prime}]$ and some induced subgraph $K$ of $[G^{\prime}]$ is such that $K\cup[f^{\prime}]$ is an atom that contains $[f^{\prime}]$ and has an $[f^{\prime}]$-preserving forbidden minor. Let $S=[C^{-1}]$, $e=[f^{\prime}]$, $G_{1}=[H^{-1}]$, and $G_{2}=K$. By Lemma 32, $H^{-1}$ is one of the graphs in Figures 20(a) - 20(m) and $C^{-1}$ is the subgraph induced by the vertex set $\{w_{i}\}$. There are two cases.
@@
Since $h$ is contained in $G^{\prime}$, it must be contained in $S$. Combining this with the fact that $e$ is retained in $[G\cup f]$ shows that $H^{-1}$ is one of the graphs in Figures 20(b), 20(f), 20(g), 20(h), 20(k), 20(l), or 20(m). There are $9$ subcases.
@@
Subcase 2: $H^{-1}$ is one of the graphs in Figures 20(f), 20(k), 20(l), or 20(m).
@@
Subcase 3: $H^{-1}$ is one of the graphs in Figures 20(g) or 20(h), $uw_{1}$ is an edge, and $h$ is $w_{2}w_{3}$ or $w_{2}w_{4}$.
@@
Subcase 4: $H^{-1}$ is one of the graphs in Figures 20(g) or 20(h), $uw_{1}$ is an edge, $h$ is $w_{3}w_{4}$, and $f^{\prime}=w_{1}w_{3}$.
Let $u$ be the endpoint of $f$ not contained in $C^{-1}$. We will show that $uw_{1}$ is reducing with respect to $(G,f)$. Let $[G\cup f]_{1}$ be the minor obtained by contracting $uw_{1}$. Observe that $f$ is retained in $[G\cup f]_{1}$, and so $uw_{1}$ is not of Type (3). We show that $uw_{1}$ is not of Type (1) or (2) by identifying some atom of $[G\cup f]_{1}$ that contains $[f]_{1}$ and has an $[f]_{1}$-preserving forbidden minor. It suffices to show that some atom of $G^{\prime}\cup\{w_{1}w_{3},w_{1}w_{4}\}$ contains $w_{1}w_{4}$ and has a $w_{1}w_{4}$-preserving forbidden minor.
@@
Subcase 5: $H^{-1}$ is one of the graphs in Figures 20(g) or 20(h), $uw_{1}$ is an edge, $h$ is $w_{3}w_{4}$, and $f^{\prime}=w_{1}w_{4}$.
This case is similar to Subcase 4. We will show that $uw_{1}$ is reducing with respect to $(G,f)$. Let $[G\cup f]_{1}$ be the minor obtained by contracting $uw_{1}$. Observe that $f$ is retained in $[G\cup f]_{1}$, and so $uw_{1}$ is not of Type (3). We show that $uw_{1}$ is not of Type (1) or (2) by identifying some atom of $[G\cup f]_{1}$ that contains $[f]_{1}$ and has an $[f]_{1}$-preserving forbidden minor. It suffices to show that some atom of $G^{\prime}\cup\{w_{1}w_{3},w_{1}w_{4}\}$ contains $w_{1}w_{4}$ and has a $w_{1}w_{4}$-preserving forbidden minor.
@@
Let $x$ be any vertex of $H^{-1}\setminus C^{-1}$ other that an endpoint of $f$. We will show that $xw_{1}$ is reducing with respect to $(G,f)$. Observe that the minor $[G\cup f]_{1}$ obtained by contracting $xw_{1}$ is $f$-retaining, and so $xw_{1}$ is not of Type (3). We show that $xw_{1}$ is not of Type (1) or (2) by identifying some atom of $[G\cup f]_{1}$ that contains $[f]_{1}$ and has an $[f]_{1}$-preserving forbidden minor.
Since $f^{\prime}$ satisfies (a), some atom $J$ of $G^{\prime}\cup f^{\prime}$ contains $f^{\prime}$ and has an $f^{\prime}$-preserving forbidden minor. We will show that either $J$ contains $w_{3}$ or $w_{2}w_{3}$ is reducing with respect to $(G^{\prime},f^{\prime})$. In the latter case, Subcase 6 applies. If $J$ does not contain $w_{3}$, then note that the minor $[G\cup f^{\prime}]_{2}$ obtained by contracting $w_{2}w_{3}$ is $f^{\prime}$-retaining, so it $w_{2}w_{3}$ is not of Type (3). Also, some CMS $E$ contained in $J$ is such that all paths in $G^{\prime}\cup f^{\prime}$ between $w_{3}$ and any vertex of $J\setminus E$ contain some vertex of $E$. Hence, $J$ is weakly retained in $[G^{\prime}\cup f^{\prime}]_{2}$. Therefore, some atom of $[G^{\prime}\cup f^{\prime}]_{2}$ contains $[J]_{2}$, and consequently contains $[f^{\prime}]_{2}$ and has an $[f^{\prime}]_{2}$-preserving forbidden minor. This shows that $w_{2}w_{3}$ is not of Type (1) or (2), and so it is reducing with respect to $(G^{\prime},f^{\prime})$.
Next, since $J$ contains $w_{3}$, $J$ is some induced subgraph $G_{2}$ of $[G^{\prime}]_{1}$ with $f^{\prime}$ as $[f^{\prime}]_{1}$. Let $G_{1}=[C^{-1}]_{1}\cup[u]_{1}$, $S=G_{1}\cap G_{2}$, and $e=[w_{1}w_{3}]_{1}$. Then, $G_{1}\cup G_{2}$ contains $[f]_{1}$. Also, since $G_{2}$ has an $[f^{\prime}]_{1}$-preserving forbidden minor, it is easy to see that $G_{1}\cup G_{2}$ has an $[f]_{1}$-preserving forbidden minor. Furthermore, since $G_{1}$ and $G_{2}$ are atoms and $e$ is a nonedge, Lemma 66 shows that $G_{1}\cup G_{2}$ is an atom. Furthermore, $G_{1}\cup G_{2}$ is contained in some atom of $[G\cup f]_{1}$ since it is an induced subgraph of $[G\cup f]_{1}$, and this atom has the desired properties.
@@
Finally, by the above arguments, some induced subgraph $I$ of $[G^{\prime}]_{1}$ is such that $I\cup[w_{1}w_{3}]_{1}$ is an atom that contains $[w_{1}]_{1}$, $[w_{3}]_{1}$, and $[w_{4}]_{1}$ and has a $[w_{1}w_{4}]_{1}$-preserving forbidden minor. Let $G_{1}=[C^{-1}]_{1}\cup[u]_{1}$, $G_{2}=I$, $S=G_{1}\cap G_{2}$, and $e=[w_{1}w_{3}]_{1}$. Similar to the argument in Subcase 8, we can use Lemma 66 to show that some atom of $[G\cup f]_{1}$ that contains $[f]_{1}$ and has an $[f]_{1}$-preserving forbidden minor.
@@
B.3   Proof of Lemmas for Proposition 9
We present proofs of Lemmas 35 and 36 in Section 5.1. Lemma 35 is proved using Lemma 56, also below.
Lemma 56 (Vertices in atom containing $f^{\prime}$).
Consider a minimal graph-nonedge pair $(G,f)$ with a top-level expanded $f$-separating CMS $C^{-1}$ whose expanded $f$-component is $H^{-1}$, and let $G^{\prime}=G\setminus(H^{-1}\setminus C^{-1})$. For any nonedge $f^{\prime}$ of $C^{-1}$, if some atom $J$ of $G^{\prime}\cup f^{\prime}$ contains $f^{\prime}$ and has an $f^{\prime}$-preserving forbidden minor and the pair $(G^{\prime},f^{\prime})$ has no reducing edge, then
  1. (i)
    $J$ contains $w_{1}$, $w_{3}$, and $w_{4}$ if $C^{-1}$ is the graph in Figure 22(b) and
  2. (ii)
    $J$ contains $C^{-1}$ if $C^{-1}$ is the graph in Figure 22(d).
@@ B.3   Proof of Lemmas for Proposition 9
Assume wlog that $f^{\prime}=w_{1}w_{3}$. Then, $J$ contains $w_{1}$ and $w_{3}$. Assume $C^{-1}$ is the graph in Figure 22(b). If $J$ does not contain $w_{4}$, then Lemma 34 states that the neighborhood of $w_{4}$ in $G^{\prime}$ is $\{w_{1},w_{3}\}$. However, Lemma 25 Statement (2) can be used to show that $w_{4}$ neighbors some vertex of $G^{\prime}$ not in $\{w_{1},w_{3}\}$. This is a contradiction, and thus $J$ contains $w_{4}$. A similar argument yields the same contradiction if $C^{-1}$ is the graph in Figure 22(d). ∎
Proof of Lemma 35.
@@
By Lemma 32, the expanded $f$-component $H^{-1}$ of $C^{-1}$ is one of the graphs in Figures 20(a) - 20(m), where $C^{-1}$ is the subgraph induced by $\{w_{i}\}$. Furthermore, if $H^{-1}$ is one of the graphs in Figures 20(c), 20(d), 20(e), 20(i), 20(k), or 20(l), then at least one dashed line-segment is a nonedge. For each case of $H^{-1}$ in Figures 20(i) - 20(m), we will show that $(G,f)$ has a reducing edge, which contradicts the minimality of this pair.
Case 1: $H^{-1}$ is one of the graphs Figures 20(i) or 20(j).
@@
We prove the lemma in the case where $H^{-1}$ is the graph Figure 20(i). The proof in the case where $H^{-1}$ is the graph Figure 20(j) is almost identical, so we omit it. Let $G^{\prime}=G\setminus(H^{-1}\setminus C^{-1})$. By Lemma 33, some atom $J$ of $G^{\prime}\cup w_{1}w_{3}$ contains $w_{1}w_{3}$ and has a $w_{1}w_{3}$-preserving forbidden minor and the pair $(G^{\prime},w_{1}w_{3})$ has no reducing edge. Also, Lemma 56 shows that $J$ contains $w_{4}$. If $J$ contains $w_{2}$, then let $x$ be any vertex of $H^{-1}\setminus C^{-1}$ that neighbors $w_{3}$. We will show that $xw_{3}$ is reducing with respect to $(G,f)$. Note that $f$ is retained in the minor $[G\cup f]$ obtained by contracting $xw_{3}$, and so this edge is not of Type (3). To see that $xw_{3}$ is not of Type (1) or (2), observe that $J$ is some induced subgraph $G_{2}$ of $[G^{\prime}]$ with $w_{1}w_{3}$ as $[w_{1}x]$. Hence, $G_{2}$ is an atom that contains $S=[C^{-1}]$ and has a $[w_{1}x]$-preserving forbidden minor. Let $G_{1}$ be the $\{[u],[w_{1}],[w_{3}]\}$-component of $[H^{-1}]$ that contains $S$. Then, $G_{1}$ is an atom and $S=G_{1}\cap G_{2}$. Since $S$ contains the nonedge $[w_{2}w_{4}]$, Lemma 66 shows that $G_{1}\cup G_{2}$ is an atom. Lastly, $G_{1}\cup G_{2}$ clearly has an $[f]$-preserving forbidden minor. Therefore, the atom of $[G\cup f]$ that contains $G_{1}\cup G_{2}$ contains $[f]$ and has an $[f]$-preserving forbidden minor.
@@
Finally, observe that $K$ is some induced subgraph $G_{1}$ of $[G^{\prime}]$ with $w_{2}w_{4}$ as $[xw_{4}]$. Let $G_{2}=[H^{-1}]$ and observe that $S=G_{1}\cap G_{2}=[C^{-1}]$ contains the nonedge $[w_{1}w_{3}]$. Since $G_{1}$ and $G_{2}$ are atoms, Lemma 66 shows that $G_{1}\cup G_{2}$ is an atom. Observe that $G_{1}\cup G_{2}$ has an $[f]$-preserving forbidden minor. These facts show that the atom of $[G\cup f]$ that contains $G_{1}\cup G_{2}$ contains $[f]$ and has an $[f]$-preserving forbidden minor. Therefore, $xw_{2}$ is not of Type (1) or (2), as desired.
Case 2: $H^{-1}$ is one of the graphs in Figures 20(k), 20(l), or 20(m).
@@
Lemma 36 follows immediately from Lemmas 57-59, below.